OpenStudy (anonymous):

Lesson 13: Functions (I will give a medal will post question in a min )

OpenStudy (anonymous):

2. Geoff planted dahlias in his garden. Dahlias have bulbs that divide and reproduce underground. In the first year, Geoff’s garden produced 4 bulbs. In the second year, it produced 8 bulbs, and in the third year, it produced 16 bulbs. If this pattern continues, how many bulbs should Geoff expect in the tenth year? (1 point) 2,048 bulbs 1,024 bulbs 40 bulbs 32 bulbs 3. What are the next three terms of the sequence 3, 6, 9, 12, …? (1 point) 15, 18, 20 15, 19, 21 15, 18, 21 14, 18, 21

OpenStudy (anonymous):

15,18,21

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

This is a Geometric progression with a = 4,r =2,So nth term = a*r^(n-1) =4*2^9 = 2048

OpenStudy (anonymous):

whats the next one?

OpenStudy (anonymous):

@Seran 4. A car traveling at 46 mi/h slows to a speed of 23 mi/h in 5 seconds. It maintains that speed for 5 seconds and then slows to a stop after 5 more seconds. Which graph shows the car’s speed over time? (1 point) hold on im gonna uploasd pic in a min

OpenStudy (anonymous):

it is b

OpenStudy (anonymous):

CAN I ASK A COUPLE MORE @Seran

OpenStudy (anonymous):

No other graphs goes to zero simply eliminate others

OpenStudy (anonymous):

yeh u r welcome

OpenStudy (anonymous):

5.Which statement describes the speed of the remote-control car over time? (1 point) A.The speed of the car decreases from 4 mi/h to 2 mi/h in the first 3 seconds, increases to 5 mi/h in the next 2 seconds, and then remains at 5 mi/h for the last 5 seconds. B.The speed of the car increases from 4 mi/h to 2 mi/h in the first 3 seconds, decreases to 5 mi/h in the next 2 seconds, and then remains at 5 mi/h for the last 5 seconds. C.The speed of the car decreases from 4 mi/h to 2 mi/h in the first 3 seconds, increases to 6 mi/h in the next second, and then remains at 6 mi/h for the last 6 seconds. D.The speed of the car decreases from 4 mi/h to 2 mi/h in the first 3 seconds, increases to 5 mi/h in the next 5 seconds, and then remains at 5 mi/h for the last 2 seconds.

OpenStudy (anonymous):

I think (C)

OpenStudy (anonymous):

It is D

OpenStudy (anonymous):

6. Which terms correctly complete the table? Time (seconds) 0 1 2 3 4 5 6 7 8 9 Distance (feet) ▒ 30 35 40 ▒ 50 55 60 65 ▒ (1 point) 25, 45, 70 25, 48, 73 20, 45, 70 25, 45, 73

OpenStudy (anonymous):

25,45,70

OpenStudy (anonymous):

Is these stuff r 4 u r home work?

OpenStudy (anonymous):

7. Which is a table of values for y = x – 6? (1 point) x y –5 1 –8 –14 –7 –13 x y –5 –11 –8 –2 –7 –13 x y –5 –11 –8 –14 –7 –13 x y –5 1 –8 –2 –7 –1

OpenStudy (anonymous):

4th set

OpenStudy (anonymous):

sry 3rd set

OpenStudy (anonymous):

8. The data in the table illustrate a linear function. x –3 0 3 6 y –6 –2 2 6 What is the slope of the linear function? Which graph represents the data? (1 point)

OpenStudy (anonymous):

@Seran

OpenStudy (anonymous):

1st one gradient = y axis difference/x axis difference = 2-(-2)/3-0 = 4/3

OpenStudy (anonymous):

9. Which hill described in the table is the steepest? Explain. Street Horizontal Distance (ft) Vertical Rise of Street (ft) Dixie Hill 80 40 Bell Hill 80 20 Liberty Hill 80 60 (1 point) A.Bell Hill; it rises 1 foot for every 4 feet of horizontal travel. B.Dixie Hill; it rises 2 feet for every 1 foot of horizontal travel. C.Liberty Hill; it rises 4 feet for every 3 feet of horizontal travel. D.Liberty Hill; it rises three fourths foot for every 1 foot of horizontal travel.

OpenStudy (anonymous):

@Seran

OpenStudy (anonymous):

ans A

OpenStudy (anonymous):

seems like u r completing all u r home works with my help is it so?

OpenStudy (anonymous):

yah and i love you 4 it and very thankful 4 it

OpenStudy (anonymous):

@Seran